# Peter Reinhard Hansen§    and   Chen Tong

### v1

@@ Abstract
Keywords: Fractional Moments, Moment-Generating Function.
@@ 1 Introduction
Moments of random variables, including conditional moments, are central to much of economics, finance, and econometrics. In this paper, we introduce a new method for obtaining moments of random variables with a well-defined moment-generating function (MGF). The proposed method is general, computationally efficient, and can provide solutions in settings where standard approaches are insufficient.
@@
It is well known that integer moments of a random variable $X$, with MGF $M_{X}(s)$, are given by $\mathbb{E}[X^{k}]=M_{X}^{(k)}(0)$, where $M_{X}^{(k)}(s)$ denotes the $k$-th derivative of the MGF. Non-integer moments and fractional absolute moments, $\mathbb{E}|X^{r}|$, $r\in\mathbb{R},$ have more complicated expressions, typically integrals that involve derivatives of the MGF or derivatives of the characteristic function (CF). [^1] Kawata, (1972) provides an expression for $\mathbb{E}|X|^{r}$ that involves several derivatives of the CF and Laue, (1980) provides an expression based on fractional derivatives, see also Wolfe, (1975), Samko et al., (1993), Matsui and Pawlas, (2016), and Tomovski et al., (2022). For fractional moments of non-negative variables, similar expressions were derived in Cressie et al., (1981), Cressie and Borkent, (1986), Jones, (1987), and Schürger, (2002). In this paper, we derive novel integral expressions for computing fractional moments, including fractional absolute moments, and fractional central moments. Unlike existing expressions, the new expressions do not involve derivatives (of the MGF or the CF). Thus, the new expressions are particularly useful when derivatives are prohibitively difficult to obtain. This is often the case in dynamic models, where the conditional MGF is defined from recursive non-linear expressions. The new method is labelled CMGF because it relies on a complex extension of the MGF, and provides expressions for complex moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{C}$. Although complex moments are not commonly used in econometrics, this generalization can be included in our results without much adaptation. Complex moments are widely applied in fields such as statistical mechanics, number theory, and quantum physics. [^2] In part, because complex moments are related to Mellin transforms. The new expressions can also be used for establishing new identities, such as the expression for the reciprocal Gamma function, which we infer from the absolute moment of Gaussian random variables.
@@
The new CMGF expressions involve integrals, and while these cannot be evaluated analytically in many situations, they facilitate new ways to compute moments numerically, which are fast and accurate. We demonstrate this with the normal-inverse Gaussian distribution, an we also illustrate the CMGF method with three well-motivated applications. First, for the Heston-Nandi GARCH (HNG) model, we obtain moments of cumulative returns. Second, we compute the moments of realized volatilities that are given by the Heterogeneous Autoregressive Gamma (HARG) model. Third, within the Autoregressive Poisson Model (APM), we compute the conditional moment for the averages number of volatility jumps in the S&P 500 index. In these applications, we are not aware of alternative methods that can compete with the CMGF method.
@@ 2 A New Method for Computing Moments (CMGF)
nests both the standard moment-generating function, $\psi(s,0)=M_{X}(s)$, and the characteristic function, $\psi(0,t)=\varphi_{X}(t)\equiv\mathbb{E}[e^{itX}]$, as special cases. The complex valued MGF, $M_{X}(z)$, is somewhat similar to the Laplace transform when $X$ is a continuous random variable. [^3] The Laplace transform of the density function, $f_{X}(x)$, is $\int_{0}^{\infty}e^{-zx}f_{X}(x)\mathrm{d}x$, and it differs from $M_{X}(z)$ in terms of the domain of integration and the sign of the exponent. Moreover, $M_{X}(z)$ is obviously also applicable to discrete distributions and distributions that are mixtures of continuous and discrete distributions.
@@
The result obviously covers regular and central absolute moments by setting $\xi=0$ for regular moments and $\xi=\mathbb{E}X$ for central moments. Note that Theorem 1 provides a continuum of expressions, one for each $s\in\mathcal{N}_{+}$. It is our experience that it makes little difference which $s\in\mathcal{N}_{+}$ is used to evaluate the moment. Although analytical integration may be impractical, (1) provides a new and alternative method to evaluate moments numerically. For real moments, $r\in\mathbb{R}$, it is our experience that it is computationally advantageous to use (2) rather than (1).
@@
$\frac{1}{\Gamma(\frac{r}{2}+1)}=\frac{2^{r/2}}{\pi}\int_{-\infty}^{\infty}\frac{e^{z^{2}/2}}{z^{r+1}}\mathrm{d}t,\qquad r>-1,$ (3)
@@
$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{+\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t.$ (5)
@@ 2.1 Example: Moments of NIG Distribution
Figure 2: Absolute moments of the standardized NIG distribution with $(\xi,\chi)=(1/2,-1/3)$ (blue lines) and $(\xi,\chi)=(1/8,-1/16)$ (red lines). These are shown in the left panels, where dots are represent the known moments for $r=2$ and $r=4$. The middle panel is a snippet of the left panel which can be compare to simulation-based moments for 15 values of $r$. For each $r$ we computed $\mathbb{E}|X|^{r}$ 100 times as represented by the x-crosses. Each of these simulation-based moments are based on $N=$1,000,000 draws of NIG distribution. The right panel shows how many decimal places are accurate with a one standard deviation simulation error.
@@
$r$ CMGF Integate with density Simulations $(N=10^{6})$
@@
The CMGF method is not only far more accurate in this example, it is also much faster than the simulation-based approach. This is shown in Table 1 where we report the computation time obtained with a Julia implementation. In this case the density is readily available, and we also compute the moment by evaluating $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$ using the same numerical integration method used to evaluate the integral in Theorem 1. The CMGF method is between 5 and 25 times faster than the standard approach, and the CMGF method is more than 10,000 times faster than averaging (and generating) $N=10^{6}$ independent draws from the NIG distribution. We have also compared the methods using a standard normal distribution, for which the true moment is known to be, $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for all $r>-1$. For the Gaussian case the CMGF is also thousands of times faster than simulation methods using 1 million draws and far more accurate. For $|r|<1$, the CMGF is also faster and more accurate than integration with the density, even after exploiting the symmetry of the density about zero. With this modification for higher moments, density integration is about twice as fast as CMGF for moments larger than one.
@@ 4 Conclusion
Future research could explore further extensions of this method to other models with the MGF in closed form. An interesting extension is to apply the CMGF method to multivariate settings, allowing for the computation of moments in multivariate distributions, including cross-moments that capture dependencies.
@@ Appendix A Appendix of Proofs
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=\frac{\Gamma(r+1)}{z^{r+1}}.$
@@
When $z\in\mathbb{R}^{+}$ (such that $z=s$, $t=0$) the result follows from the definition of the gamma function, $\Gamma(r+1)=\int_{0}^{\infty}y^{r}e^{-y}\mathrm{d}y=\int_{0}^{+\infty}s^{r}x^{r}e^{-sx}s\mathrm{d}x$, where we used integration by substitution with $y=sx$.
For the general situation, $z=s+it\in\mathbb{C}^{+}$, we consider
$\lim_{K\rightarrow\infty}\int_{0}^{K}x^{r}e^{-zx}\mathrm{d}x=\frac{1}{z^{r+1}}\int_{0}^{sK+itK}y^{r}e^{-y}\mathrm{d}y,$
where we used the substitution: $y=zx\in\mathbb{C}$. The function in the last integral, $f(y)\equiv y^{r}e^{-y}$, has a pole at $y=0$ if $r<0$, and is a multi-valued function when $r$ is a non-integer. We will adopt the principal value by restricting the angle of $y=\lambda e^{i\theta}$ to $\theta\in(-\pi,\pi]$, and construct a closed contour integral that avoids the pole at $y=0$. The contour is illustrated in Figure 7.
Figure 7: The closed contour for the integral, $\oint y^{r}e^{-y}\mathrm{d}y$.
The function, $f(y)$, is analytic over the closed contour for $\lambda>0$, (i.e. it does not include any singularity). Therefore, by Cauchy’s Integral Theorem, we have
The integral we seek is therefore given by,
This completes the proof. ∎
@@
For $x<0$, $z=s+it\in\mathbb{C}^{+}$, and $r+1\in\mathbb{C}^{+}$ we have
@@
For $x=0$, the identity holds for $r\in\mathbb{C}^{+}$.
@@
because the last term is zero for $r$ with ${\rm Re}\left(r\right)>0$.
@@
We will use a closed contour integral in the complex plane with two part. The first part is the line segment along the real axis, from $-R$ to $R$, and the second part is the semicircular arc, $C_{R}=\{Re^{i\theta},\theta\in[0,\pi]\}$. The integral over the line segment, $\int_{-R}^{+R}f(\tau)\mathrm{d}\tau$, is the object of interest, whereas the semicircle defines an auxiliary integral that is used to close the contour, such that we can apply Cauchy’s Integral Theorem. The first reason for the change of variable, $\tau=-t$, is to place the pole of $f(\zeta)$ outside the closed contour for all $R>0$ (the pole is at $\zeta=-is$), and by Cauchy’s Integral Theorem, we have $0=\int_{-R}^{R}f(\zeta)\mathrm{d}\zeta+\int_{C_{R}}f(\zeta)\mathrm{d}\zeta$. For the integral along the arc, we have the expression,
@@
Proof of Lemma 1. We first focus on the case where $x\geq 0$. With $\mathrm{Re}[z]=s>0$, from Lemma A.1, we have the identity
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=$ $\frac{\Gamma(r+1)}{z^{r+1}},\qquad\text{for all}\quad r+1\in\mathbb{C}^{+}.$
@@
$x^{r}=\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\frac{\Gamma(r+1)}{z^{r+1}}e^{zx}\mathrm{d}z,\quad x>0,$
@@
$|x|^{r}=x^{r}$ $=\frac{\Gamma(r+1)}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}z,$
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,$ (A.1)
which is well-defined for all $r+1\in\mathbb{C}^{+}$ (because $x>0$). For $x=0$ it is well defined for all $r\in\mathbb{C}^{+}$. This finishes the proof of the third formula in Lemma 1. We proceed to get
@@
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\quad x\geq 0$ (A.2)
where the second term is zero from Lemma A.2.
Then, we focus on the case that $x<0$. From the identity A.1, we directly have
$|x|^{r}=\left(-x\right)^{r}$ $=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t,\quad x<0$
Similarly, we proceed to get
$|x|^{r}$ $=\left(-x\right)^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t+\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t,\quad x<0,$ (A.3)
where the second term is zero from Lemma A.2.
Combine the two formulas (A.2) and (A.3), we have the first formula in Lemma 1
$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\quad x\in\mathbb{R}.$
The second formula in Lemma 1, from expression (A.1), we have
$x^{k}$ $=\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{k+1}}\mathrm{d}t,\quad x>0,$
$x^{k}$ $=\left(-1\right)^{k}\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{k+1}}\mathrm{d}t,\quad x<0,$
by using the similar steps in the proof of first formula in Lemma 1, we have
$x^{k}=\frac{k!}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+\left(-1\right)^{k}e^{-zx}}{z^{k+1}}\mathrm{d}t,$
where we use $\Gamma(k+1)=k!$ for $k\in\mathbb{N}$. This completes the proof. $\square$
@@
for $r+1\in\mathbb{C}^{+}$ if $\Pr(X=0)=0$ and $r\in\mathbb{C}^{+}$ if $\Pr(X=0)>0$. The same line of arguments applies for continuous distribution, given by

### v2

@@ Abstract
Keywords: Moments, Fractional Moments, Moment-Generating Function.
@@ 1 Introduction
Moments of random variables, including conditional moments, are central to much of economics, finance, and econometrics. In this paper, we introduce a new method for obtaining moments of random variables with a well-defined moment-generating function (MGF). The proposed method is general, computationally efficient, and can offer solutions to problems where standard methods are inadequate.
@@
It is well known that the $k$-th integer moment of a random variable $X$, with MGF, $M_{X}(s)=\mathbb{E}[e^{sX}]$, is given by $\mathbb{E}[X^{k}]=M_{X}^{(k)}(0)$, where $M_{X}^{(k)}(s)$ is the $k$-th derivative of $M_{X}(s)$. Non-integer moments and fractional absolute moments, $\mathbb{E}|X^{r}|$, $r\in\mathbb{R},$ have more complicated expressions, typically integrals that involve derivatives of the MGF or derivatives of the characteristic function (CF). [^1] Kawata, (1972) provides an expression for $\mathbb{E}|X|^{r}$ that involves several derivatives of the CF and Laue, (1980) provides an expression based on fractional derivatives, see also Wolfe, (1975), Samko et al., (1993), Matsui and Pawlas, (2016), and Tomovski et al., (2022). For fractional moments of non-negative variables, similar expressions were derived in Cressie et al., (1981), Cressie and Borkent, (1986), Jones, (1987), and Schürger, (2002). In this paper, we derive novel integral expressions for computing fractional moments, including fractional absolute moments, and fractional central moments. Unlike existing expressions, the new expressions do not involve derivatives (of the MGF or the CF). Thus, the new expressions are particularly useful when derivatives are prohibitively difficult to obtain. This is often the case in dynamic models, where the conditional MGF is defined from recursive non-linear expressions. The new method is labelled CMGF because it relies on a complex extension of the MGF, and provides expressions for complex moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{C}$. Although complex moments are not commonly used in econometrics, this generalization can be included with minimal adaptation.  [^2] Complex moments are commonly used other fields, including statistical mechanics, number theory, and quantum physics. In part, because complex moments are related to Mellin transforms. The new expressions can also be used for establishing new identities, such as the expression for the reciprocal Gamma function, which we infer from the absolute moment of a Gaussian random variable.
@@
The new CMGF expressions involve integrals, and while these cannot be evaluated analytically in most cases, they facilitate new ways to compute moments numerically, which are fast and accurate. We demonstrate this with the normal-inverse Gaussian distribution. We also show that the CMGF method offers fast and accurate moments in dynamic models, where we are not aware of good alternative methods. Specifically, we use the CMGF method to compute moments of cumulative returns in the Heston-Nandi GARCH (HNG) model, to compute moments of realized volatilities in the Heterogeneous Autoregressive Gamma (HARG) model, and to compute conditional moments in the Autoregressive Poisson Model.
@@ 2 A New Method for Computing Moments (CMGF)
nests both the standard moment-generating function, $\psi(s,0)=M_{X}(s)$, and the characteristic function, $\psi(0,t)=\varphi_{X}(t)\equiv\mathbb{E}[e^{itX}]$, as special cases. The complex valued MGF, $M_{X}(z)$, is the bilateral Laplace–Stieltjes transform (apart from the sign of $z$), and for continuous random variables it is similar to the Laplace transform.  [^3] The Laplace transform of the density function, $f_{X}(x)$, is $\int_{0}^{\infty}e^{-zx}f_{X}(x)\mathrm{d}x$, and it differs from $M_{X}(z)$ in terms of the domain of integration and the sign of the exponent. Moreover, $M_{X}(z)$ is obviously also applicable to discrete distributions and distributions that are mixtures of continuous and discrete distributions.
@@
The result obviously covers regular and central absolute moments by setting $\xi=0$ for regular moments and $\xi=\mathbb{E}X$ for central moments. Note that the expression remains invariant to the specific value of $s$, so long as $s\in\mathcal{N}_{+}$. In our experience, the choice of $s$ has no practical impact on moment evaluation. Although analytical integration may often be impractical, (1) offers a novel method to evaluate moments numerically. For real moments, $r\in\mathbb{R}$, it is our experience that it is computationally advantageous to use (2) rather than (1).
@@
$\frac{1}{\Gamma(\frac{r}{2}+1)}=\frac{2^{r/2}}{\pi}\int_{-\infty}^{+\infty}\frac{e^{z^{2}/2}}{z^{r+1}}\mathrm{d}t,\qquad r>-1,$ (3)
@@
$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t.$ (5)
@@ 2.1 Example: Moments of NIG Distribution
Figure 2: Absolute moments of the standardized NIG distribution with $(\xi,\chi)=(1/2,-1/3)$ (blue lines) and $(\xi,\chi)=(1/8,-1/16)$ (red lines). These are shown in the left panels, where dots are represent the known moments for $r=2$ and $r=4$. The middle panel is a snippet of the left panel which can be compared to simulation-based moments for 15 values of $r$. For each $r$ we computed $\mathbb{E}|X|^{r}$ 100 times as represented by the x-crosses. Each of these simulation-based moments are based on $N=$1,000,000 draws of NIG distribution. The right panel shows how many decimal places are accurate with a one standard deviation simulation error.
@@
$r$ CMGF Integrate with density Simulations $(N=10^{6})$
@@
The CMGF method is not only far more accurate in this example, it is also much faster than the simulation-based approach. This is shown in Table 1 where we report the computation time obtained with a Julia implementation. In this case the density is readily available, and we also compute the moment by evaluating $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$ using the same numerical integration method used to evaluate the integral in Theorem 1. The CMGF method is between 5 and 25 times faster than the standard approach, and the CMGF method is more than 10,000 times faster than averaging (and generating) $N=10^{6}$ independent draws from the NIG distribution. We have also compared the methods using a standard normal distribution, for which the true moment is known to be, $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for all $r>-1$. For the Gaussian case the CMGF is also thousands of times faster than simulation methods using 1 million draws and far more accurate. For $|r|<1$, the CMGF is also faster and more accurate than numerical integration based on the density (even if we take advantage of the density being symmetric about zero). For moments larger than one, integration with the density is about twice as fast as CMGF, after taking advantage of the symmetric density.
A multivariate extension is beyond the scope of this paper. However, some simple cross-moments, such as $\mathbb{E}[X_{1}X_{2}]$ and $\mathbb{E}[X_{1}X_{2}^{2}]$, are easy to obtain from univariate results. Let $(X_{1},X_{2})$ have MGF, $M_{X}(t_{1},t_{2})$, such that $M_{X_{1}+X_{2}}(t)=M_{X}(t,t)$, $M_{X_{1}-X_{2}}(t)=M_{X}(t,-t)$, $M_{X_{1}}(t)=M_{X}(t,0)$, and $M_{X_{2}}(t)=M_{X}(0,t)$, from which we obtain $\mathbb{E}[(X_{1}+X_{2})^{k}]$, $\mathbb{E}[(X_{1}-X_{2})^{k}]$, etc. by Theorem 3, and the two cross moments are given by $\mathbb{E}[X_{1}X_{2}]=\frac{1}{2}\left(\mathbb{E}[(X_{1}+X_{2})^{2}]-\mathbb{E}[X_{1}^{2}]\mathbb{E}[X_{2}^{2}]\right)$ and $\mathbb{E}[X_{1}X_{2}^{2}]=\frac{1}{6}\left(\mathbb{E}[(X_{1}-X_{2})^{3}]+\mathbb{E}[(X_{1}+X_{2})^{3}]-2\mathbb{E}[X_{1}^{3}]\right)$, respectively.
@@ 4 Conclusion
Future research could explore further extensions of this method to other models with closed-form MGFs. An interesting extension is to apply the CMGF method to multivariate distributions, allowing for the computation of moments in multivariate distributions, including cross-moments that capture dependencies, beyond the simplest cross-moments, $\mathbb{E}[X_{1}X_{2}]$, discussed in Section 2.
@@ Appendix A Appendix of Proofs
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=\frac{\Gamma(r+1)}{z^{r+1}}.$ (A.1)
@@
Consider the case where $z$ is real and positive, i.e. $z=s>0$. From the definition of the gamma function and the substitution, $y\mapsto sx$, we have $\Gamma(r+1)=\int_{0}^{\infty}y^{r}e^{-y}\mathrm{d}y=\int_{0}^{\infty}s^{r}x^{r}e^{-sx}s\mathrm{d}x$, which proves the identity for any $z\in\mathbb{R}^{+}$. Both sides of (A.1) are holomorphic functions for $z\in\mathbb{C}^{+}$, and since they agree on $z\in\mathbb{R}^{+}$, it follows by the identity theorem that they agree for all $z\in\mathbb{C}^{+}$. [^9] We thank Christian Berg for suggesting this simplification of the proof. A different, direct proof for all $z\in\mathbb{C}^{+}$, is presented in the Supplementary Material. ∎
@@
For $x<0$, $\operatorname{Re}(z)>0$, and $\operatorname{Re}(r)>-1$ we have
@@
For $x=0$, the identity holds for $\operatorname{Re}(r)>0$.
@@
because the last term is zero for ${\rm Re}\left(r\right)>0$.
@@
We will use a closed contour integral in the complex plane with two part. The first part is the line segment along the real axis, from $-R$ to $R$, and the second part is the semicircular arc, $C_{R}=\{Re^{i\theta},\theta\in[0,\pi]\}$. The integral over the line segment, $\int_{-R}^{+R}f(\tau)\mathrm{d}\tau$, is the object of interest, whereas the semicircle defines an auxiliary integral that is used to close the contour, such that we can apply the Cauchy Integral Theorem. The first reason for the change of variable, $\tau=-t$, is to place the pole of $f(\zeta)$ outside the closed contour for all $R>0$ (the pole is at $\zeta=-is$), such that $0=\int_{-R}^{R}f(\zeta)\mathrm{d}\zeta+\int_{C_{R}}f(\zeta)\mathrm{d}\zeta$ by the Cauchy Integral Theorem. For the integral along the arc, we have the expression,
@@
Proof of Lemma 1. Consider first the case $x\geq 0$. Since $\mathrm{Re}[z]=s>0$, we have the following identity from Lemma A.1,
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=$ $\frac{\Gamma(r+1)}{z^{r+1}},\qquad\text{for all}\quad\mathrm{Re}[r]>-1.$
@@
$x^{r}=\frac{1}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{\Gamma(r+1)}{z^{r+1}}e^{zx}\mathrm{d}z,\quad x>0,$
@@
$|x|^{r}=x^{r}=\frac{\Gamma(r+1)}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}z=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,$ (A.2)
which is well-defined for all $\mathrm{Re}[r]>-1$ (because $x>0$). For $x=0$ the expression is well-defined for all $\mathrm{Re}[r]>0$. This completes the proof of the third expression in Lemma 1. We proceed with
@@
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\qquad\quad x\geq 0,$ (A.3)
where the second term is zero by Lemma A.2.
For the case $x<0$, we have the following identity directly from (A.2),
$|x|^{r}=\left(-x\right)^{r}$ $=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad x<0.$
Adding $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ to the right-hand side, which equals zero Lemma A.2, we arrive at
$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for }x<0.$ (A.4)
Combining (A.3) and (A.4) proves the first expression in Lemma 1.
From (A.2), we also have
$x^{k}$ $=\qquad\quad\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{+zx}}{z^{k+1}}\mathrm{d}t,\quad x>0,$
$x^{k}$ $=\left(-1\right)^{k}\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{k+1}}\mathrm{d}t,\quad x<0.$
This, and $\Gamma(k+1)=k!$ for $k\in\mathbb{N}$, proves the second expression in Lemma 1. This completes the proof. $\square$
@@
for $r+1\in\mathbb{C}^{+}$ if $\Pr(X=0)=0$ and $r\in\mathbb{C}^{+}$ if $\Pr(X=0)>0$. For a continuous distribution, the same line of arguments gives us,